Hello,
Please help me solve this problem, thanks in advance.
5/2x + 1/4x=11/4+x

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

11/4x=11/4+x
11/4x-4/4x=11/4
7/4x=11/4
x=11/4*4/7
x=44/28
x=1 4/7

Answer by stanbon(75887)

5/2x + 1/4x=11/4+x
------
What is the denominator of the 11?
Is the fraction 11/4 or 11/(4+x)?
Cheers,
Stan H.